I have a question that staes: given f(x)=8/x+6 and g(x)=6/x+3, find (f-g)(x).
Answer by stanbon(75887)

given f(x)=8/x+6 and g(x)=6/x+3, find (f-g)(x).
--------
(f-g)x means f(x)-g(x)
--------
(f-g)x = 8/(x+6) - 6/(x+3)
=[8x+24-6x-36]/(x+6)(x+3)
=[2x-12]/(x+6)(x+3)
=[2(x-6)]/[(x+6)(x+3)]
===========
Cheers,
Stan H.